The radius of a silver bar is 30 π‘π‘š. What force is required to stretch the bar by



12.7% of its length where Young’s modulus is π‘Œ = 7.25 Γ— 1010 π‘π‘šβˆ’2

1
Expert's answer
2021-12-06T09:48:43-0500

According to Hooke's law


"F=EA\cdot\frac{\Delta l}{l}=E\cdot\pi r^2\cdot\frac{0.127l_0}{l_0}=0.127\cdot E\cdot\pi r^2="


"=0.127\cdot7.25\cdot10^{10}\cdot 3.14\cdot0.3^2=2.6\cdot10^9\ (N)" . Answer